# Asynchronous Message-Passing and Zeroth-Order Optimization Based Distributed Learning with a Use-Case in Resource Allocation in Communication Networks

### v2

@@ Abstract
Distributed learning and adaptation have received significant interest and found wide-ranging applications in machine learning and signal processing. While various approaches, such as shared-memory optimization, multi-task learning, and consensus-based learning (e.g., federated learning and learning over graphs), focus on optimizing either local rewards or a global reward, there remains a need for further exploration of their interconnections. This paper specifically focuses on a scenario where agents collaborate towards a common task (i.e., optimizing a global reward equal to aggregated local rewards) while effectively having distinct individual tasks (i.e., optimizing individual local parameters in a local reward). Each agent’s actions can potentially impact other agents’ performance through interactions. Notably, each agent has access to only its local zeroth-order oracle (i.e., reward function value) and shares scalar values, rather than gradient vectors, with other agents, leading to communication bandwidth efficiency and agent privacy. Agents employ zeroth-order optimization to update their parameters, and the asynchronous message-passing between them is subject to bounded but possibly random communication delays. This paper presents theoretical convergence analyses and establishes a convergence rate for nonconvex problems. Furthermore, it addresses the relevant use-case of deep learning-based resource allocation in communication networks and conducts numerical experiments in which agents, acting as transmitters, collaboratively train their individual policies to maximize a global reward, e.g., a sum of data rates.
@@ I-A Distributed learning over message-passing architectures
The recent proliferation of edge devices with computational power, data generation and processing, has motivated research on distributed learning and adaptation. In distributed learning, edge devices, as distributed agents, aim to cooperatively learn a common task or multiple individual tasks. This collaboration occurs within a message-passing architecture, which poses challenges due to heterogeneity in local reward functions/data, processing capabilities, and potential communication delays or failures.
Asynchronous distributed optimization and learning techniques in the literature address this collaboration and consider various message-passing approaches as follows: Shared-memory approach: All agents have access to a common shared memory [1, 2, 3, 4, 5]. Each agent reads the required information from the shared memory and updates the decision variable, either in full or in part, by writing to the shared memory. This model operates asynchronously as the reading and writing processes occur asynchronously. Local-memory approach: Each agent possesses a local (private) memory and asynchronously updates a specific part of the decision variable, i.e., the local parameters [5, 6, 4, 7]. After each update, the agent broadcasts the updated local parameter to other users who require it for their own updates. Due to random communication delays or dropouts, the information received from other agents and stored in the local memories may be outdated. The local-memory approach can be considered, for instance, in multitask learning approaches [8] where each agent optimizes its local parameters based on the sum of its local reward function and a regularizer that captures the relation of the local parameters with other agents’ local parameters. In games and multi-agent reinforcement learning, this approach can be considered where each agent optimizes its local reward function while capturing its dependence on other players through interaction [9, 10]. Consensus-based approach: Each agent also updates its local parameters asynchronously. However, unlike the previous approaches, agents in this approach aim to achieve parameter consensus. Consensus can be achieved through frequent averaging facilitated by a central entity, as seen in federated learning settings [11, 12, 13], or through frequent averaging of local parameters among neighboring agents, as seen in learning over graph settings [14, 15, 16, 17, 18, 19]. In this approach, all agents strive to reach common optimal parameters optimized over a global reward function, such as the sum of local rewards.
@@
In this setting, each agent’s actions and decisions can potentially impact the performance of other agents, necessitating consideration of interactions during local optimization, similar to the local-memory approach. Furthermore, agents optimize their local parameters not only based on their own local rewards but also with respect to a global function (e.g., sum of local rewards) that reflects the common task, similar to the consensus-based approach.
To the best of our knowledge, this is the first theoretical convergence study addressing such a scenario, where agents asynchronously update their local parameters using zeroth-order oracles by querying only local reward function values. Employing zeroth-order optimization with zeroth-order oracles rather than first-order oracles (i.e., function gradients) in each agent brings multiple benefits. While first- or higher-order oracles may not be available or expensive to calculate in some settings (e.g., simulation-based optimization or bandit optimization [19]), zeroth-order oracles can be available in a more computationally- and memory-efficient manner [20]. Moreover, zeroth-order oracles significantly reduce the communication bandwidth requirements in the proposed distributed learning as it helps agents share only scalars. This feature, sharing scalars unlike sharing local parameters or gradients in federated learning or learning over graph settings, ensures privacy-preserving cooperation [21] (cf. Section V for more details). The relevant problem of RA is examined as a use-case within this framework.
@@ I-B Resource allocation in communication networks
Distributed training facilitates online training of RA policies meaning that the policies can be trained simultaneously while they are being utilized. This offers several advantages over centralized (offline) training: (1) Online training enables the capture of real-time channel and user state distributions, resulting in more accurate policy training compared to offline training, which may suffer from a mismatch between the training sample distribution and the actual sample distribution [24]. (2) Non-stationary channels can lead to distribution drifts, which are caused by the dynamic nature of communication systems. Factors such as transmitter and receiver dynamic locations, varying environmental conditions (e.g., varying traffic patterns in vehicle-to-vehicle communication), and shifting user demand patterns can contribute to distribution drifts [38, 28]. Distributed training provides a means to address these distribution mismatches and drifts through (online) policy retraining [28]. (3) In practical scenarios, it may be infeasible to have a server with sufficient communication and computational capacity to perform centralized training. Distributed training offers a viable alternative in such cases, as it does not rely on a central server and can distribute the training process across multiple agents [39]. Online (distributed) training can be established by employing zeroth-order or higher-order optimization approaches.
Zeroth-order optimization for RA policy training (also called model-free training) is considered in this paper. During the proposed zeroth-order training, agents approximate utility function gradients by measuring their utility values, e.g., data rates, rather than calculating function gradients [24]. This approach can benefit from the advantages of zeroth-order optimization mentioned in the previous subsection. Moreover, it captures the full behavior of the communication system, including nonidealities such as nonlinearities in the modulator and demodulator or antenna setup, which are not typically captured by higher-order approaches due to assumed simplifications in the utility functions, e.g., simplifications made in data rate formulation. The contributions are listed as follows:
@@
    A novel distributed learning approach is proposed, where agents share a common task (hence they optimize the sum of local rewards), while each agent has also an individual task (hence agents have local (unique) parameters). Agents impact each others’ performance through interactions. They update their local parameters using zeroth-order optimization methods. During the training, agents share scalar quantities, which saves communication bandwidth and preserves their privacy.
@@ II System Model
As an example, the optimization problem in (1) deals with a scenario where there are $m$ agents, possibly with different tasks $\bm{\theta}_{i}\neq\bm{\theta}_{j},i\neq j$. However, they also share a common task $f(\bm{\theta})$. Therefore, each agent $\ell$ seeks to optimize its local parameters $\bm{\theta}_{\ell}$ over the sum of all local reward functions. This is different from multi-task or game scenarios where each agent $\ell$ addresses the problem given as $\operatorname*{minimize}_{\bm{\theta}_{\ell}}f_{\ell}(\bm{\theta})$, with $f_{\ell}(\bm{\theta})=g_{\ell}(\bm{\theta}_{\ell})+R(\bm{\theta})$ consisting of the local reward $g_{\ell}$ and a regularizer capturing the relation between tasks $\bm{\theta}_{j}$ in multi-task scenarios [8] or $f_{\ell}(\bm{\theta})$ representing the local reward influenced by other players in the game [9]. In addition, it is different from federated learning [11, 12, 13] or learning over graphs [14, 15, 16, 17, 18, 19] where agents seek a consensus in $\operatorname*{minimize}_{\bm{\theta}=(\bm{\theta}_{1},\cdots,\bm{\theta}_{m})\in\mathbb{R}^{n}}\sum_{\ell=1}^{m}g_{\ell}(\bm{\theta}_{\ell})$ by including the constraint $\bm{\theta}_{\ell}=\bm{\theta}_{j},\ell\neq j$. As an illustrative use-case, we focus on the relevant problem of resource allocation (RA) in wireless communication systems.
@@
Consider the problem of RA where agents are represented as $m$ transmitters in the communication network, serving $n\leq m$ receivers. The set of transmitters and receivers are denoted as $\mathcal{T}$ and $\mathcal{R}$, respectively. In the case where $m=n$, the communication network is considered ad-hoc, i.e., each transmitter serves only one dedicated receiver. In the case where $m>n$, the receivers can be considered as base stations, where multiple transmitters have uplink communication with one base station, possibly causing interference to other base stations. We refer to $r(i)\in\mathcal{R}$ as the receiver dedicated to transmitter $i\in\mathcal{T}$. The communication channel between the $i$th transmitter and the $j$th receiver is considered as a random variable, denoted by $h_{ji}\in\mathbb{C}$. The matrix $\bm{H}\in\mathbb{C}^{|\mathcal{R}|\times|\mathcal{T}|}$, with ||$\cdot$|| ⋅ | representing the cardinality of the set, contains all the channels in the communication network, with the channel $h_{ji}$ at the $j$th row and the $i$th column of the matrix.
The RA problem aims to find optimal resources (e.g., transmit power) for each user (as an agent) such that the global reward (e.g., sum of user data rates for communication) is maximized. This is done by performing the optimization (1) with $f_{\ell}(\bm{\theta})=\mathbb{E}_{\bm{\xi}=\bm{H}}{\mathopen{}\left[r^{\pi}_{\ell}(\bm{\pi}(\bm{\theta},\bm{H}),\bm{H})\mid{{}}\right]\mathclose{}}$ where $r^{\pi}_{\ell}:\mathbb{R}^{m}\times\mathbb{C}^{|\mathcal{R}|\times|\mathcal{T}|}\rightarrow\mathbb{R}$ denotes agent $\ell$’s reward or so-called utility, e.g., its data rate. The vector $\bm{\pi}(\bm{\theta},\bm{H})\coloneqq(\pi_{1}(\bm{\theta}_{1},\bm{H}),\cdots,\pi_{m}(\bm{\theta}_{m},\bm{H}))\in\mathbb{R}^{m}$ contains individual RA policies $\pi_{i}:\mathbb{R}^{n_{i}}\times\mathbb{C}^{|\mathcal{R}|\times|\mathcal{T}|}\rightarrow\mathbb{R}$, e.g., deep neural networks (DNNs). These policies map their input to communication resources, e.g., the transmit power of the corresponding agents. Each policy $\pi_{i}$ is parametrized by the policy parameter $\bm{\theta}_{i}\in\mathbb{R}^{n_{i}}$. Due to interference between agents through cross-talk channels $h_{ji},\leavevmode\nobreak\ j\neq i$, the utility of each agent, i.e., $r^{\pi}_{\ell}$, depends on all the policies $\bm{\pi}$. Therefore, cooperative learning is required. Otherwise, the utilities are decoupled as $r^{\pi}_{\ell}(\pi_{\ell}(\bm{\theta}_{\ell},\bm{H}),h_{r(\ell)\ell})$ which can be addressed by non-cooperative learning.
@@ III Proposed Learning Setup
In this section, we present the setup that governs the local updates and communication among the agents. Frequently used notations are listed in Table I in the Appendix.
@@ III-A Agent Activity Time
Let $t\in\mathbb{N}$ represent the global time. The set of time instants at which agent $i$ updates its local parameters is defined as $T^{i}_{u}$. Furthermore, each agent $i$ queries its zeroth-order oracle (i.e., measures its reward function value) $F_{i}(\bm{\theta}^{t},\bm{\xi})$ at certain time instants, which define the set $T^{i}_{q}$. At any time, each agent is free to decide whether to update its parameter or keep the current parameter, as well as whether to query its oracle or not. Consequently, we take the following assumption:
@@ III-B Message-passing Architecture
The exchange of information among agents depends on the structure of the local reward functions $f_{i}$ for all $i\in[m]$. To facilitate the presentation of this exchange, we introduce neighbor sets for each agent $i\in[m]$ as follows:
@@
(a) Message-passing architecture among the agents. The quantities $\bar{r}_{j}^{t}$ and $\mathcal{T}_{i}^{j,t}$ are defined in (4) and (5), respectively. Depending on its local reward function, each agent defines the sets of agents $\mathcal{N}^{i}_{rx}$ and $\mathcal{N}^{i}_{tx}$ with which it communicates and saves necessary variables in its local memory. Due to asynchrony and random communication delays, the memory contents may be outdated.
(b) Update and query time of agent $1$. The axis ticks correspond to global time instants $t\in\mathbb{N}$. Agents asynchronously query their reward functions and update their local parameters.
@@ IV Zeroth-order optimization
The technique is called randomized smoothing, as the approximation function $f^{\mu}$ is smooth (cf. I for definition) even if the function $f$ is not [41, lem. 2]. The computation of the gradient ${\nabla}f^{\mu}$ in Item 2 constitutes the key computational task for each agent in the proposed learning approach, as elaborated in the following section. Importantly, this gradient derivation solely relies on querying the zeroth-order oracle, i.e., the reward function value, using perturbed and unperturbed decision variables. It is noteworthy that throughout the learning process, the smooth approximation function $f^{\mu}$ is optimized, rather than the original function $f$. However, as indicated by Item 3, it becomes evident that by selecting a sufficiently small constant $\mu$, the two functions can be made arbitrarily close to each other.

### v3

@@ Abstract
Distributed learning and adaptation have received significant interest and found wide-ranging applications in machine learning and signal processing. While various approaches, such as shared-memory optimization, multi-task learning, and consensus-based learning (e.g., federated learning and learning over graphs), focus on optimizing either local costs or a global cost, there remains a need for further exploration of their interconnections. This paper specifically focuses on a scenario where agents collaborate towards a common task (i.e., optimizing a global cost equal to aggregated local costs) while effectively having distinct individual tasks (i.e., optimizing individual local parameters in a local cost). Each agent’s actions can potentially impact other agents’ performance through interactions. Notably, each agent has access to only its local zeroth-order oracle (i.e., cost function value) and shares scalar values, rather than gradient vectors, with other agents, leading to communication bandwidth efficiency and agent privacy. Agents employ zeroth-order optimization to update their parameters, and the asynchronous message-passing between them is subject to bounded but possibly random communication delays. This paper presents theoretical convergence analyses and establishes a convergence rate for nonconvex problems. Furthermore, it addresses the relevant use-case of deep learning-based resource allocation in communication networks and conducts numerical experiments in which agents, acting as transmitters, collaboratively train their individual policies to maximize a global reward, e.g., a sum of data rates.
@@ I-A Distributed learning over message-passing architectures
The recent proliferation of edge devices with computational power, along with increased data generation and processing, has motivated research on distributed learning and adaptation. In distributed learning, edge devices, as distributed agents, aim to cooperatively learn a common task or multiple individual tasks. This collaboration occurs within a message-passing architecture, which poses challenges due to heterogeneity in local cost functions/data, processing capabilities, and potential communication delays or failures.
Asynchronous distributed optimization and learning techniques in the literature address this collaboration and consider various message-passing approaches as follows: Shared-memory approach: All agents have access to a common shared memory [1, 2, 3, 4, 5]. Each agent reads the required information from the shared memory and updates the decision variable, either in full or in part, by writing to the shared memory. This model operates asynchronously as the reading and writing processes occur asynchronously. Local-memory approach: Each agent possesses a local (private) memory and asynchronously updates a specific part of the decision variable, i.e., the local parameters [5, 6, 4, 7]. After each update, the agent may broadcast the updated local parameter to other users who require it for their own updates. Due to random communication delays or dropouts, the information received from other agents and stored in the local memories may be outdated. The local-memory approach can also be considered, for instance, in multitask learning approaches [8] where each agent optimizes its local parameters based on the sum of its local cost function and a regularizer that captures the relation of the local parameters with other agents’ local parameters. In games and multi-agent reinforcement learning, this approach can be considered where each agent optimizes its local cost function while capturing its dependence on other players through interaction [9, 10]. Consensus-based approach: Each agent updates its local parameters (asynchronously). However, unlike the previous approaches, agents in this approach aim to achieve parameter consensus. Consensus can be achieved through frequent averaging facilitated by a central entity, as seen in federated learning settings [11, 12, 13], or through frequent averaging of local parameters among neighboring agents, as seen in learning over graph settings [14, 15, 16, 17, 18, 19]. In this approach, all agents strive to reach common optimal parameters optimized over a global cost function, such as the sum of local costs.
@@
In this setting, each agent’s actions and decisions can potentially impact the performance of other agents, necessitating consideration of interactions during local optimization, similar to the local-memory approach. Furthermore, agents optimize their local parameters not only based on their own local costs but also with respect to a global function (e.g., sum of local costs) that reflects the common task, similar to the consensus-based approach.
To the best of our knowledge, this is the first theoretical convergence study addressing such a scenario, where agents asynchronously update their local parameters using zeroth-order oracles by querying only local cost function values. Employing zeroth-order optimization with zeroth-order oracles rather than first-order oracles (i.e., function gradients) in each agent brings multiple benefits. While first- or higher-order oracles may not be available or expensive to calculate in some settings (e.g., simulation-based optimization or bandit optimization [19]), zeroth-order oracles can be available in a more computationally- and memory-efficient manner [20]. Moreover, zeroth-order oracles significantly reduce the communication bandwidth requirements in the proposed distributed learning as it helps agents share only scalars. This feature, sharing scalars unlike sharing local parameters or gradients in federated learning or learning over graph settings, ensures privacy-preserving cooperation as well [21] (refer to Section V for more details). The relevant problem of resource allocation (RA) has been examined as a use-case within this framework.
@@ I-B Resource allocation in communication networks
Distributed training facilitates online training of RA policies meaning that the policies can be trained simultaneously while they are being utilized. This offers several advantages over centralized (offline) training: (1) Online training enables the capture of real-time channel and user state distributions, resulting in more accurate policy training compared to offline training, which may suffer from a mismatch between the training sample distribution and the actual sample distribution [24]. (2) Non-stationary channels can lead to distribution drifts, which are caused by the dynamic nature of communication systems. Factors such as transmitter and receiver dynamic locations, varying environmental conditions (e.g., varying traffic patterns in vehicle-to-vehicle communication), and shifting user demand patterns can contribute to distribution drifts [38, 28]. Distributed training provides a means to address these distribution mismatches and drifts through (online) policy retraining [28]. (3) In practical scenarios, it may be infeasible to have a server with sufficient communication and computational capacity to perform centralized training. Distributed training offers a viable alternative in such cases, as it does not rely on a central server and can distribute the training process across multiple agents [39].
Online (distributed) training can be established by employing zeroth-order or higher-order optimization approaches. Zeroth-order optimization for RA policy training (also called model-free training) is considered in this paper. During the proposed zeroth-order training, agents approximate utility function gradients by measuring their utility values, e.g., data rates, rather than calculating function gradients [24]. This approach can benefit from the advantages of zeroth-order optimization mentioned in the previous subsection. Moreover, it captures the full behavior of the communication system, including nonidealities such as nonlinearities in the modulator and demodulator or antenna setup, which are not typically captured by higher-order approaches due to assumed simplifications in the utility functions, e.g., the simplifications made in data rate formulation. The contributions are listed as follows:
@@
    A novel distributed learning approach is proposed, where agents share a common task (hence they optimize the sum of local costs), while each agent has also an individual task (hence agents have local (unique) parameters). Agents impact each others’ performance through interactions. They update their local parameters using zeroth-order optimization methods. During the training, agents share scalar quantities, which saves communication bandwidth and preserves their privacy.
@@ II System Model
As an example, the optimization problem in (1) deals with a scenario where there are $m$ agents, possibly with different tasks $\bm{\theta}_{i}\neq\bm{\theta}_{j},i\neq j$. However, they also share a common task $f(\bm{\theta})$. Therefore, each agent $\ell$ seeks to optimize its local parameters $\bm{\theta}_{\ell}$ over the sum of all local cost functions. This is different from multi-task or game scenarios where each agent $\ell$ addresses the problem given as $\operatorname*{minimize}_{\bm{\theta}_{\ell}}f_{\ell}(\bm{\theta})$, with $f_{\ell}(\bm{\theta})=g_{\ell}(\bm{\theta}_{\ell})+R(\bm{\theta})$ consisting of the local cost $g_{\ell}$ and a regularizer $R$ capturing the relation between tasks $\bm{\theta}_{j}$ in multi-task scenarios [8] or $f_{\ell}(\bm{\theta})$ representing the local cost influenced by other players in the game [9]. In addition, it is different from federated learning [11, 12, 13] or learning over graphs [14, 15, 16, 17, 18, 19] where agents seek a consensus in $\operatorname*{minimize}_{\bm{\theta}=(\bm{\theta}_{1},\cdots,\bm{\theta}_{m})\in\mathbb{R}^{n}}\sum_{\ell=1}^{m}g_{\ell}(\bm{\theta}_{\ell})$ by including the constraint $\bm{\theta}_{\ell}=\bm{\theta}_{j},\ell\neq j$. As an illustrative use-case, we focus on the relevant problem of RA in wireless communication systems.
@@
Consider the problem of RA where agents are represented as $m$ transmitters in the communication network, each serving a single receiver, with a total number of receivers $n\leq m$. The set of transmitters and receivers are denoted as $\mathcal{T}$ and $\mathcal{R}$, respectively. In the case where $m=n$, the communication network is considered ad-hoc, i.e., each transmitter serves only one dedicated receiver. In the case where $m>n$, the receivers can be considered as base stations, where multiple transmitters have uplink communication with one base station, possibly causing interference to other base stations. We refer to $r(i)\in\mathcal{R}$ as the receiver dedicated to transmitter $i\in\mathcal{T}$. The communication channel between the $i$th transmitter and the $j$th receiver is considered as a random variable, denoted by $h_{ji}\in\mathbb{C}$. The matrix $\bm{H}\in\mathbb{C}^{|\mathcal{R}|\times|\mathcal{T}|}$, with ||$\cdot$|| ⋅ | representing the cardinality of the set, contains all the channels in the communication network, with the channel $h_{ji}$ at the $j$th row and the $i$th column of the matrix.
The RA problem aims to find optimal resources (e.g., transmit power) for each user (as an agent) such that the global reward (e.g., sum of user data rates for communication) is maximized. This is done by performing the optimization (1) with $f_{\ell}(\bm{\theta})=-\mathbb{E}_{\bm{\xi}=\bm{H}}{\mathopen{}\left[r^{\pi}_{\ell}(\bm{\pi}(\bm{\theta},\bm{H}),\bm{H})\mid{{}}\right]\mathclose{}}$ where $r^{\pi}_{\ell}:\mathbb{R}^{m}\times\mathbb{C}^{|\mathcal{R}|\times|\mathcal{T}|}\rightarrow\mathbb{R}$ denotes agent $\ell$’s reward or so-called utility, e.g., its data rate. The vector $\bm{\pi}(\bm{\theta},\bm{H})\coloneqq(\pi_{1}(\bm{\theta}_{1},\bm{H}),\cdots,\pi_{m}(\bm{\theta}_{m},\bm{H}))\in\mathbb{R}^{m}$ contains individual RA policies $\pi_{i}:\mathbb{R}^{n_{i}}\times\mathbb{C}^{|\mathcal{R}|\times|\mathcal{T}|}\rightarrow\mathbb{R}$, e.g., DNNs. These policies map their input to communication resources, e.g., the transmit power of the corresponding agent. Each policy $\pi_{i}$ is parametrized by the policy parameter $\bm{\theta}_{i}\in\mathbb{R}^{n_{i}}$. Due to interference between agents through cross-talk channels $h_{ji},\leavevmode\nobreak\ j\neq i$, the utility of each agent, i.e., $r^{\pi}_{\ell}$, depends on all the policies $\bm{\pi}$. Therefore, cooperative learning is required. Otherwise, the utilities are decoupled as $r^{\pi}_{\ell}(\pi_{\ell}(\bm{\theta}_{\ell},h_{r(\ell)\ell}),h_{r(\ell)\ell})$ and the RA problem can be addressed trivially or by non-cooperative learning.
@@ III Proposed Learning Setup
In this section, we present the setup that governs the local updates and communication among the agents. The frequently used notations are listed in Table I in the Appendix.
@@ III-A Agent Activity Time
Let $t\in\mathbb{N}$ represent the global time. The set of time instants at which agent $i$ updates its local parameters is defined as $T^{i}_{u}$. Furthermore, each agent $i$ queries its zeroth-order oracle (i.e., measures its cost function value) $F_{i}(\bm{\theta}^{t},\bm{\xi})$ at certain time instants, which define the set $T^{i}_{q}$. At any time, each agent is free to decide whether to update its parameter or keep the current parameter, as well as whether to query its oracle or not. Consequently, we take the following assumption:
@@ III-B Message-passing Architecture
The exchange of information among agents depends on the structure of the local cost functions $f_{i}$ for all $i\in[m]$. To facilitate the presentation of this exchange, we introduce neighbor sets for each agent $i\in[m]$ as follows:
@@
(a) Message-passing architecture among the agents. The quantities $\bar{r}_{j}^{t}$ and $\mathcal{T}_{i}^{j,t}$ are defined in (4) and (5), respectively. Depending on its local cost function, each agent defines the sets of agents $\mathcal{N}^{i}_{rx}$ and $\mathcal{N}^{i}_{tx}$ with which it communicates and saves necessary variables in its local memory. Due to asynchrony and random communication delays, the memory contents may be outdated.
(b) Update and query time of agent $1$. The axis ticks correspond to global time instants $t\in\mathbb{N}$. Agents asynchronously query their cost functions and update their local parameters.
@@ IV Zeroth-order optimization
The technique is called randomized smoothing, as the approximation function $f^{\mu}$ is smooth (refer to I for definition) even if the function $f$ is not [41, lem. 2]. The computation of the gradient ${\nabla}f^{\mu}$ in Item 2 constitutes the key computational task for each agent in the proposed learning approach, as elaborated in the following section. Importantly, this gradient derivation solely relies on querying the zeroth-order oracle, i.e., the cost function value, using perturbed and unperturbed decision variables. It is noteworthy that throughout the learning process, the smooth approximation function $f^{\mu}$ is optimized, rather than the original function $f$. However, as indicated by Item 3, it becomes evident that by selecting a sufficiently small constant $\mu$, the two functions can be made arbitrarily close to each other.